Commissioning Notes from 12-13 May 2008 MST

Observers: J. Hill and J. Storm on the mountain, plus D. Miller and A. Rakich from Tucson
Telescope Operator: R. Grashuis

Summary

Sky was clear almost the entire night, but an approaching cold front gives us high wind 8-16 m/sec, and bad seeing 1.5 to 4 arcsec. Guider was working decently well in 1.5 arcsec seeing at the beginning of the night, although there are still some things I don't understand. Doug and Andrew got some more calibration measurements to study the transformation between S-H zernikes and M1 zernikes. We ran out of time before closing the active optics loop with the S-H sensor. Jesper and John used the second half to measure coordinate transformations for the guide probe. We measured 6 stars around M5_0619, and we took a few standard star images near NGC6341. It is hard to make astrometric measurements in 3 arcsec seeing, but we got a decent start. We lost 25 minutes to TCS reflective memory problems on lbtmu101.

Nighttime Activities

01:55 Open dome just before sunset. Wind from SW at a steady 15 m/sec. Sky clear, but lots of smoke and haze down low. T=6degC.

02:00 D. Ashby running tests of a new interpolator algorithm.

02:45 Dave is finished. Telescope to zenith so Randy can turn down the last of the treehouse fans.

03:20 pointto WT10_365. We are working out of the wind in the NE.

03:23 repeated instances of "MCSPU error returned - ERROR : there are no tracking polynomials." Stop and restart PCS. We have a new build of PCS, but the GCS is the same as last night.

03:29 Remove SX Pointing offsets and clear active optics. Start spiral search, since we didn't see the star. Found at offset -30 60.

While adjusting the zeropoint of the pointing model.....
Input term name: IE     Input term value: -28
PCS setAuthorizedPMTerm failed; fatal error in tcsPup
Michele's trap works, it did not crash PCS.

03:45 net by-eye active optics correction +3400 Z8, with adjustPointing left M1 -10 0 to provide range. Why do we have this big left-right effect compared to two weeks ago?

Input term name: IE     Input term value: -28
PCS setAuthorizedPMTerm failed; fatal error in tcsPup
This still moves the star! Is it a spurious error?

03:55 Finished adjusting pointing model to zero the pointing offset: IE=-27 CA=--105 for offset 0 0

04:02 pointto WT10_365 gstar=0 GUIDE. Why is guide star out of focus compared to IRTC? (Jesper thinks the nominal value of 21-23 mm was calculated for an off-axis position, so the on-axis focus of the guide probe should be ~4 mm higher.)

04:06 setfocus -p 24

04:07 setGuideRate 0.4

Wavefront Sensing and Active Optics Tests

04:10 pointto BS9152 gstar=0 GUIDE. Failed to find star at 1 msec exposure. Use GCS command line to take a 1000 msec Guider exposure. This is a star suitably faint to do wavefront sensing with multi-second exposures.

04:12 pointto again, starts guiding but still some jumping around,

04:15 Adjusted SX Zglobal 0.09 to make star in focus on Guider rather than IRTC.

04:18 WFS 0001 image 5 sec (double spots) 04:19 WFS 0002 image 15 sec

May 12 21:18:58 lbtmu107 LBT_Networkserver: GCSL died unexpectedly.

Start GCSL, but it doesn't start until you send it a stop command. It sort of half starts, but doesn't register any RPC commands until Randy does netconfig stop GSCL. Then it registers the RPC commands and is running OK.

04:25 pointto BS9152 gstar=0 GUIDE.

4:25 setGuideRate 0.4

04:28 WFS image 15 sec times a few. Guide images is bouncing by +-1.5 arcsec. This star is giving ~8000 counts peak per subaperture in 15 sec.

04:30 Doug is taking WFS images (in a directory other than /tmp where mine are) in /Repository/AGW_Data/080513.

04:37 Press Collimate button.

04:38 Adjusted SX Zglobal 0.15 mm to improve focus on guider. Guiding seems better when the focus is better. Maybe GCS was measuring hotspots in the distorted pupil. Now the star is moving +-0.3 arcsec.

Why is guiding better tonight?
  • better seeing (1.5 versus 3 arcsec)
  • more steady wind
  • better focus on guider camera

04:51 Press Collimate button

05:00 setfilter 2 (oac is zero-based), Torsten thinks this is filter 3.

05:08 Press Collimate button

05:22 Press Collimate button (It's a tough job, but someone has to do it.)

Guider is doing fairly well, but it still has a tendency to be slightly left of the hotspot (like 286 instead of 292).

05:34 Press Collimate button at EL=70deg.

05:55 Guider image was bouncing around. More wind gusts? So the WFS spots are sausage shaped.

06:01 Apply Z5=+1000 nm (Doug measures -1450 Z5 on wfsc0035. Already includes multiplication by 2)

06:07 Apply Z5=-1000 nm (net zero) (Doug measures -308 nm Z5 on wfsc0038.)

06:15 Apply Z6=+1000 nm (Doug measures -104 Z5, +1480 Z6 on wfsc0041.) Sign reversal on Z6.

06:19 Apply Z6=-1000 nm (net zero) (wfsc0044 is rubbish because I moved while exposing)

06:24 Apply Z7=+1000 nm

06:34 Apply Z7=-1000 nm (net zero)

Lost guide star when I bumped into the X=+2.6 mm limit with +1000 Z8.

06:47 Apply Z8=-1000 nm (wfsc0054)

06:49 pointto, didn't capture the guide star on the first try. OK on the second try.

6:54 setGuideRate 0.4

AGw Calibration

07:05 Randy sent preset from PresetGUI. This failed to stop GCS guiding. Move to NGC6341. Stars in NGC6341 to faint for easy ID with IRTC in 5 sec.

07:18 pointto M5_0619 gstar=1 TRACK

07:22 pointto M5_0619 PA=0 TRACK IE=-32 CA=-114

takepic 0.3

07:25 setxy -x 0 -y 612.5

readGuideCam 500, 100, 10

07:29 pointo gstar 3 PA=0 TRACK setxy -3.3 517.4

no star in guider in 1.0 sec or 10 sec.

07:34 pointto gstar=3 PA=180 TRACK

IE=-36 CA=-114

07:36 guider000038 1 sec shows gstar #3 centered (256,252)

setxy -1.3 517.4

00:39 guider000039 1 sec shows gstar #3 centered (335,253)

gstar=6 PA=180

setxy -5.3 455.4

07:45 guider000040 183,152

IE=-37 CA=-112

07:47 guider000041 154,131

setxy -3.3 455.4

07:48 guider000042 (217,133)

setxy -2.3 457.4

07:49 guider000043 (246,188)

setxy -2.0 460.4

07:53 guider000044 (257,296)

setxy -1.1 459.2

07:56 guider000045 (286,265) 000046

gstar=3 PA=180

setxy -1.5 517.4

08:02 guider000047 (310,250)

gstar=3 PA=179.45 pointto_080316.log

setxy -1.0 510.7

08:06 guider000048 (301,35)

setxy -1.0 517.3

08:07 guider000049 (295,242)

gstar=4 PA=0 pointto_080929.log

IE+ -33.5 CA=-111

setxy -1.0 510.7

08:13 guider000050 (39,219)

setxy +6.5 511.8

08:15 guider000051 (270,257)

setxy 7.1 511.8

08:17 guider000052 (285,254)

gstar=5 PA=0 pointto_082204.log

setxy +33.5 506.0

08:24 guider000053 (225,173)

setxy +35.3 508.4

08:25 guider000054 (275,253)

gstar=2 PA=0

setxy +65.6 +566.1

08:29 guider000055 (285, 134)

setxy 65.6 569.7

08:30 guider00056 (271,250) guider000057 (275,245)

IE=-34 CA=-110.5

08:33 guider000058 (281,247)

gstar=1 PA=0

setxy -12.3 542.0

08:35 guider000059 (249,257)

08:42 adjustPointing left M1 -15 0

08:46 "Remove Corrections" for Pointing offset. Star moves but numbers on PSF GUI don't change. Stopping GUI doesn't help.

08:47 netconfig stop PSFL, and restart. Numbers in Pointing Offsets still don't clear, but now collimate is sending garbage to PMCL. At same time Randy is having problems with ECS. It seems like there is a TCS meltdown underway. PCS is running OK on lbtmu102. Both PSFL and ECS are running on lbtmu101. We restart ECS on lbtmu103 and it works OK. Now we stop and restart MCS on lbtmu104 and it works. Stop PSFL and move it to lbtmu103 and it is OK. So apparently lbtmu101 has become completely disconnected from the rest of reflective memory.

Lost ~25 minutes to Issue #1344.

09:13 Clear SX M1 active optics, and recollimate with adjustPointing left M1 -20 0 Z7=+500, Z8=+3400 (near +X limit) CA=-98 IE=-35

Active X=+2.902 Y=+0.427 Z=0 RX=4.6 RY=-31.2 at EL=52.9 Pointing X=-0.931 Y=0 Z=0 RX=0 RY=+20 (for a -20 adjust pointing) These should go into global offsets.

09:25 start a focus sequence for AGw, although the seeing fluctuations are pretty ugly. Maybe later........

09:31 pointto M5_0619 PA=0 again since rotator wasn't tracking, pointto_093052.log. In the meantime, Jesper has installed a new coordinate transformation.

Beware of 1 msec exposures in variable seeing!

09:35 guider000062 24pix FWHM at Focus=21

09:37 guider000063 28 pix FWHM at Focus=25

09:38 guide000064 is saturated at Focus=29, so this star is too bright.

09:39 guider000065 is also saturated at Focus=21, no info here.

setxy 0.0 512.5 (semi-random offset)

09:41 guider000067 star ? at (53, 277)

Jesper updates again

09:43 guider000068 star 4 at (54,284)

M5_0619 star offsets in relative arcsec (RA/DEC Guide Star - RA/DEC Target Star). Second pair of numbers are the offsets in radians Third pair of numbers are the offsets in mm computed by "pointto". Fourth pair of numbers are the offsets in mm using the transformation that Jesper calculated the following day. Last pair are the positions we measured on the sky.

GS dRA dDEC dRArad dDECrad X mm Y mm fitX fitY skyX skyY
star 4 -1.63 +169.29 -7.9025e-6 8.2074e-4 -0.978 510.92 2.46 513.1 7.1 511.8
star 1 -20.46 +119.69 -9.9193e-5 5.8027e-4 -12.276 540.686 -9.59 541.8 -12.3 542.0
star 5 +56.34 +177.46 2.7314e-4 8.6035e-4 33.804 506.024 36.6 509.5 35.3 508.4
star 2 +109.78 +76.52 5.3223e-4 3.7098e-4 65.868 566.588 66.0 569.8 65.4 569.7
star 3 +5.50 -158.91 2.6665e-5 -7.7042e-4 3.300 707.846 0.0 705.9
star 6 +8.69 -258.78 4.2130e-5 -1.2546e-3 5.214 767.768 -0.1 764.6

10:06 restart oacserver with new transformation

10:12 Preset to NGC6341 star D, can't see in 4 sec at H on IRTC, but it was out of focus because the preset GUI doesn't collimate M1

getfocus 40,

setxy 0 612.5,

10:16 guider000070 5 sec

setxy 2 612.5

setxy 5 612.5

setxy 0 612.5 (center) guider000074

10:21 CA=-80 guider000075

IE=-25

10:22 IE=-40 guider000077 (famous ghost?

IE=-60

IE=-80 Not a ghost but a pupil, and those images in the last 10 min were all bogus because we didn't collimate at this elevation.

10:25 IE=-40 guider000080 (428,165)

10:27 CA=-90 guider000083 (344,325)

IE=-33

10:29 Preset to NGC6341 star G, now we see 2 medium stars and 5 faint stars in 1 sec, but seeing is horrible (PA=0)

10:33 setfocus 25 guider000086

10:34 setfocus 0 guider000088 definitely pupils about 100 pixels diameter

10:37 setfocus 48 guider000089

Images are really blurry at all focus positions, but Polaris monitor is reporting above 4 arcsec (some combination of windshake and seeing.

10:39 setfocus 30 guider000090 36 pixels FWHM (1.8 arcsec) 1 sec exposure

10:41 guider000091 5 sec exposure

setxy 0 512, IRTC is only barely detecting these stars in 5 sec.

10:45 guider000092 is 5 sec at xy 0 512

10:47 setxy 0 612.5 guider000093 is 5 sec back on axis

10:48 setxy 0 552 guider000094 is 4 sec

10:48 setxy 50 500 guider000095

10:49 setxy -50 500 guider000096

10:50 setxy 0 612.5 guider000097 is back on axis

10:54 above was with oacfilter 3, now switch to oacfilter 2 guider000098

10:55 now switch to oacfilter 1 guider000099

10:56 now switch to oacfilter 0 guider000100

10:57 now switch to oacfilter 4 (empty?) guider000101

Bug: if you ask for oacfilter 5, then it reports 1, and after that everything is messed up and you don;t get the filter you ask for.

11:00 home -m 16 to home the filter wheel.

11:11 Hand oacserver over to Jesper for fiddling

Randy and John did the start azimuth axis in wind experiment. The axis restarted impressively well quartering a 10 m/sec wind (Closed Issue #847.

11:27 pointto M5_0619 again at PA-0

setxy -x -4.679005 -y 514.333988 for star 4

CA=-98 IE=-38

11:29 guider000102 no star

setxy -x 4.679005 -y 514.333988 for star 4 (-X)

11:31 guider000103 found star at 213,335 but faint (clouds? sky is mostly clear but some cirrus in NW)

setxy -x -14.493712 -y 543.533951 for star 1

11:40 guider000105 (147,294)

11:46 restart oacserver

11:48 restart oacserver (back to the 4AM version)

11:49 restart oacserver (this transformation didn't work)

11:51 restart oacserver

12:12 Close Dome T=-1degC. Wind 8m/sec from south.

Copied the acquisition images with off-axis stars into /Repository/AGW_Data/080513g/

Plan for Tomorrow Night

Partly cloudy. Chance of snow showers with chance of thunderstorms. Probability of measurable precipitation 40 percent. Northwest wind 23 gusting to 35 mph. Low 26. We will continue tonight's activities if the forecast is wrong.

-- JohnHill - 13 May 2008
Topic revision: r7 - 15 May 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback